OpenStudy (anonymous):

Use an angle sum identity to verify the identity cos 2 theta = 2 cos^2 theta - 1

OpenStudy (anonymous):

what is angle sum identity...I can figure out which one it is on this list

OpenStudy (anonymous):

So, we need to verify: \[cos(2\theta)=2cos^2(\theta)-1\] A angle sum formula is one which allows us to calculate angles in the form of \(\theta=a+b\) Where \(a\) and \(b\) are portions of the angle. For \(\{cos\}\), we have: \[cos(a+b)=cos(a)cos(b)-sin(a)sin(b)\] So, we can say that \(cos(2\theta)=cos(\theta+\theta)\). Let \(a=\theta\), and \(b=\theta\); we can derive the following equation: \(cos(\theta+\theta)=cos(2\theta)\) And \(cos(\theta+\theta)=cos(\theta)cos(\theta)-sin(\theta)sin(\theta)=cos^2(\theta)-sin^2(\theta)\) So therefore, \(cos(2\theta)=cos^2(\theta)-sin^2(\theta)\). We can simplify this a little further by remembering that \(cos^2(\theta)+sin^2(\theta)=1\) and observing that \(sin^2(\theta)=1-cos^2(\theta)\) Therefore, we can simplify the equation further to: \[\eqalign{cos(2\theta)&=cos^2(\theta)-sin^2(\theta) \\ &=cos^2(\theta)-[1-cos^2(\theta)] \\ &=cos^2(\theta)-1+cos^2(\theta) \\ &=2cos^2(\theta)-1 \\}\]